OpenStudy (anonymous):

Which value for x makes the open sentence true? 5 + 4 • x = 32 + x2 A. 5 B. 3 C. 4 D. 2

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

\[5+4 \times x=32+x^2\] solve for x, do you know how?

OpenStudy (anonymous):

bunki it should be c

OpenStudy (anonymous):

\[5+4x-32-x^2=0\]rearrange\[-x^2+4x-27=0\]\[x^2-4x+27=0\]\[SOLVE!\]

OpenStudy (anonymous):

thank you all c;

OpenStudy (anonymous):

I don't think it is c though it doesn't work if you plug it in.

OpenStudy (anonymous):

\[(x-2)^2=-27+4\]\[x=2±isqrt{23}\]

OpenStudy (anonymous):

\[x=2±i \sqrt{23}\]